# New AGN diagnostic diagrams based on the [OIII] auroral line

### v1

@@ 3.4 Defining the locus of AGN-only objects
For the three diagnostic diagrams described above, we were able to trace clear demarcation lines providing a separation between the AGN-dominated regions and the regions where SFGs and AGN can overlap. To do so, we considered both the photoionization models and the observational sample distributions. We ultimately define the demarcation lines in the form:
where the values of the parameters and for each diagnostic are reported in Table 2. It is important to note that the AGN selection based on these demarcation lines is not a necessary condition for a source to be selected as an AGN, but rather a sufficient condition: objects above the demarcation line are AGN-dominated with high confidence, while objects below the demarcation line can either be SFGs or AGN.
O3Hg vs O32 | O3Hg vs Ne3O2 | O3Hg vs O33 | |
a | 0.55 | 0.48 | -1.1 |
b | -0.95 | -0.42 | 1.47 |
@@ 4 Discussion
By comparing the distribution of the observational samples with the distribution of the photoionization models, we can explain why these diagnostic diagrams can separate part of the AGN population from SFGs. In FigĀ 6 of the appendix, we plot the AGN and SFG models in the same diagnostic diagrams presented in Sec.Ā 3, but highlighting the variation of and , i.e. the two main parameters affecting the distribution of the models in these plots. We know that, since they are very local, SDSS AGN and SFG samples are populated mostly by solar metallicity and moderate-to-low ionization parameter sources, as shown also by the models occupying the same region of these samples in FigĀ 6. On the contrary, local analogs are characterized, on average, by lower metallicities and higher ionization parameters (as high-z galaxies). Indeed, in Fig.Ā 1 and Fig.Ā 2, they fill the regions that extend from the SDSS SFG towards higher [OIII]/Hand higher [OIII]/[OII]or [NeāIII]/ [OII], in close agreement with the trend of photoionization models for higher and in the same diagnostics (see Fig.Ā 6).
The reason for the presence, in particular in Fig.Ā 1 and Fig.Ā 2, of a region that is populated only by AGN, at high [OIII]/Hvalues, can be explained in the following way. Since the [OIII]line is a collisionally excited line generated from high energy levels, its intensity relative to His directly related to the temperature and, secondly, to the metallicity of the ISM and ionization parameter. The main difference between the AGN and SFGs is the SED of the incident ionizing radiation. For AGN the ionizing radiation is given by the emission of the accretion disk, while for SFGs it originates from young star clusters. The former are characterized by a harder spectrum and, therefore, on average photons are more energetic. Higher energy per ionizing photon, deposited into the ISM, results into a higher effective heating, hence higher electron temperature (hence higher [OIII]) at a given ionization parameter.
@@ 4.1 Stacking of the AGN candidates
To further demonstrate the effectiveness of the diagnostic diagrams proposed in this work we explored potential tracers of AGN activity in those high-z galaxies lying in the AGN region of the diagnostics diagrams in Fig. 1 and Fig. 2 but not previously identified as AGN based on the selection performed in Scholtz etĀ al. (2023a) and Mazzolari et al. 2024 (in prep).
@@
To model the broad wing in the H, we refit the Hand [NāII]Ā with two additional models - BLR model (by fitting one additional broad Gaussian component to Honly) and an outflow model by adding broad Gaussian components to both Hand [NāII]Ā with fixed FWHM and centroid. Using the Bayesian Information Criterion (BIC) parameter (Liddle, 2007), we found that the BLR model is strongly preferred to the narrow-only fit (). The broad component has an SNR of 10, making this a solid detection. On the contrary, the fit with the outflow model did not return any broad component in the [NāII]Ā lines.
@@
The absence of a counterpart in [OIII]strongly disfavours an outflow origin of the broad component of H. Therefore, the detection of a broad component in the Hemission line indicates the presence of an AGN BLR in the spectra of at least some of these galaxies, which is too weak to be detected in the individual spectra. This indicates that galaxies located in the upper region of our diagnostic diagrams, even if originally not classified as AGN, actually may host faint TypeI AGN. This result provides confidence that the diagnostic diagrams presented in this work are effective in identifying AGN above our proposed demarcation lines, although they cannot distinguish between AGN and SFGs below the demarcation line.
@@ 5 Conclusions
We stress that being above the demarcation line in the proposed diagrams is a sufficient but not necessary condition for an object to be identified as an AGN. Galaxies located below such demarcation lines can be powered by star formation and/or AGN.
@@ Acknowledgements
GM acknowledges useful conversations with Roberto Gilli and Marcella Brusa. FDE, JS and RM acknowledge support by the Science and Technology Facilities Council (STFC), from the ERC Advanced Grant 695671 āQUENCHā, and by the UKRI Frontier Research grant RISEandFALL. RM also acknowledges funding from a research professorship from the Royal Society. HĆ gratefully acknowledges support by the Isaac Newton Trust and by the Kavli Foundation through a Newton-Kavli Junior Fellowship.

### v2

@@ 3.4 Defining the locus of AGN-only objects
For the three diagnostic diagrams described above, we were able to trace clear demarcation lines providing a separation between the AGN-dominated regions and the regions where SFGs and AGN can overlap. To do so, we considered both the photoionization models and the observational sample distributions. The demarcation lines for the O3Hg vs O32 diagnostic diagram are defined as follows:
Similarly, for the O3Hg vs Ne3O2 diagnostic diagram, the demarcation lines are:
Finally, the demarcation line for the O3Hg vs O33 diagnostic diagram is the following:
These demarcation lines ensure that the fraction of contaminating SDSS SFG in the AGN part of the diagnostics is less than 1-2%, while none of the local analogs lie in the AGN-only part of the three diagnostics.
Such a tiny fraction of contaminants can probably be associated with sources that are not identified as AGN by the BPT diagram (used to distinguish between SDSS SFG and AGN in this work), but that host active SMBHs. The presence of AGN contaminants among BPT-selected SFG and local analogs has already been demonstrated by recent works (Svoboda etĀ al., 2019; Harish etĀ al., 2023).
Furthermore, it is important to note that the AGN selection based on these demarcation lines is not a necessary condition for a source to be selected as an AGN, but rather a sufficient condition: objects above the demarcation line are AGN-dominated with high confidence, while objects below the demarcation line can either be SFGs or AGN.
@@ 4 Discussion
By comparing the distribution of the observational samples with the distribution of the photoionization models, we can explain why these diagnostic diagrams can separate part of the AGN population from SFGs. In FigĀ 6, we plot the AGN and SFG models in the same diagnostic diagrams presented in Sec.Ā 3, but highlighting the variation of and , i.e. the two main parameters affecting the distribution of the models in these plots. We know that, since they are very local, SDSS AGN and SFG samples are populated mostly by solar metallicity and moderate-to-low ionization parameter sources, as shown also by the models occupying the same region of these samples in FigĀ 6. On the contrary, local analogs are characterized, on average, by lower metallicities and higher ionization parameters (as high-z galaxies). Indeed, in Fig.Ā 1 and Fig.Ā 2, they fill the regions that extend from the SDSS SFG towards higher [OIII]/Hand higher [OIII]/[OII]or [NeāIII]/ [OII], in close agreement with the trend of photoionization models for higher and in the same diagnostics (see Fig.Ā 6).
The reason for the presence, in particular in Fig.Ā 1 and Fig.Ā 2, of a region that is populated only by AGN, at high [OIII]/Hvalues, can be explained in the following way. Since the [OIII]line is a collisionally excited line generated from high energy levels, its intensity relative to His directly related to the temperature and, secondly, to the metallicity of the ISM and ionization parameter. The main difference between the AGN and SFGs is the SED of the incident ionizing radiation. For AGN the ionizing radiation is given by the emission of the accretion disk, while for SFGs it originates from young star clusters. The former are characterized by a harder spectrum and, therefore, on average photons are more energetic (see Fig.1 in Feltre etĀ al., 2016). Higher energy per ionizing photon, deposited into the ISM, results into a higher effective heating, hence higher electron temperature (hence higher [OIII]) at a given ionization parameter.
@@ 4.1 Stacking of the AGN candidates
To further demonstrate the effectiveness of the diagnostic diagrams proposed in this work we explored potential tracers of AGN activity in those high-z galaxies lying in the AGN region of the diagnostics diagrams in Fig. 1 and Fig. 2 but not previously identified as AGN based on the selection performed in Scholtz etĀ al. (2023a) and Mazzolari et al., in prep.
@@
To model the broad wing in the H, we refit the Hand [NāII]Ā with two additional models - BLR model (by fitting one additional broad Gaussian component to Honly) and an outflow model by adding broad Gaussian components to both Hand [NāII]Ā with fixed FWHM and centroid. Using the Bayesian Information Criterion (BIC) parameter (Liddle, 2007), we found that the BLR model is strongly preferred to the narrow-only fit (). The broad component has an S/N of 10, making this a solid detection. On the contrary, the fit with the outflow model did not return any broad component in the [NāII]Ā lines.
@@
The absence of a counterpart in (brighter) [OIII]strongly disfavours an outflow origin of the broad component of H. Therefore, the detection of a broad component in the Hemission line suggests the presence of an AGN BLR in the spectra of at least some of these galaxies, which is too weak to be detected in the individual spectra. This indicates that galaxies located in the upper region of our diagnostic diagrams, even if originally not classified as AGN, actually may host faint TypeI AGN. This result provides confidence that the diagnostic diagrams presented in this work are effective in identifying AGN above our proposed demarcation lines, although they cannot distinguish between AGN and SFGs below the demarcation line.
@@ 5 Conclusions
We stress that being above the demarcation line in the proposed diagrams is a sufficient but not necessary condition for an object to be identified as an AGN. Galaxies located below such demarcation lines can be powered by star formation and/or AGN. At the same time, the contamination from SFG is almost completely negligible in the upper part of the diagnostics, as demonstrated by the distribution of low-redshift observational samples, photoionization models. Moreover also the few high-redshift ambigous cases in the AGN-only region turn out to be AGN based on the stacking.
@@ Acknowledgements
We thank the anonymous referee for useful suggestions which improved the quality of the paper.
GM acknowledges useful conversations with Roberto Gilli, Marcella Brusa, Sandro Tacchella, William Baker, Callum Witten, Bartolomeo Trefoloni, Lola Danhaive, Ignas Juodzbalis, Amanda Stoffer, Brian Xing Jiang, and William McClaymont.
FDE, JS and RM acknowledge support by the Science and Technology Facilities Council (STFC), from the ERC Advanced Grant 695671 āQUENCHā, and by the
UKRI Frontier Research grant RISEandFALL. RM also acknowledges funding from a research professorship from the Royal Society.
HĆ gratefully acknowledges support by the Isaac Newton Trust and by the Kavli Foundation through a Newton-Kavli Junior Fellowship.